OpenStudy (anonymous):

The membership fee for joining a camping association is $45. A local campground charges members of the camping association $35 per night for a campsite and non members $40 per night for a campsite. After how many nights of camping is the total cost for members, including the membership fee, the same as the total cost for nonmembers?

OpenStudy (anonymous):

45+35x = 40x (x is the number of nights) 45 = (40 - 35)x 45 = 5x 45/5 = x x = 9 nights

OpenStudy (anonymous):

Thankyou!